Grade 12Physical Chemistry

How can three resistors of resistances 2 Ω, 3 Ω, and 6 Ω be connected to give a total resistance of
1 Ω,

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

R1 = 2 ohm
R2 = 3 ohm
R3 = 6 ohm
When R1,R2, R3 are connected in parallel we get
1/R = 1/R1 + 1/R2 + 1/R3
= 1/2 + 1/3 + 1/6
= 1 ohm.

Thanks